PG:

how many positive two-digit numbers are evenly divisible by 4

PG:

HELP ME

Alexis1415:

" A number is divisible by 4 if and only if its last two digits (ten's place and one's place) are divisible by 4." The total numbers divisible by 4 is 22. Based on this quote you can solve it by seeing what it can go into.

surjithayer:

first two digit number divisible by 4 is 12. 4)99(24 96 - 3 99-3=96 so last digit=96 sequence is 12,16,20,...,96 first term a=12 last term l=96 c.d. d=16-12=4 l=a+(n-1)d 96=12+(n-1)4 96-12=(n-1)4 84/4=n-1 n-1=21 n=21+1=22 total two digit terms divisible by 4=22